# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v8

@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
1. Average daily wage ($w^{Daily}_{j}$): The average daily wage of male factory workers in the skilled age range (aged 30-40) in the machinery sector is calculated using two official reports of the manufacturing censuses documented in the ASCT and LST. Details in the calculation steps are summarized in Online Appendix B.6.1.
2. Average annual working days ($l_{j}$): The ASCT documents the number of factories and average annual working days for each machinery sub-sector in 2 factory-size bins. I used the number of factories in each machinery sub-sector and factory size bin as the weights to calculate the weighted average of the annual working days in the entire machinery sector.
@@
Sources: Data used to calculate the average daily wage are from the Tokyo City Office (1921, pp. 726-757); Tokyo City Office (1922, pp. 732-733); Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data used to calculate the average annual working days are from the Tokyo City Office (1921, pp. 726-741) and Tokyo City Office (1922, pp. 716-725). Data used to calculate the average share of monthly ancillary wage are from Kitazawa (1924, pp. 20; 58-62) and the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674).
Following the institution described in Online Appendix B.6.2, the monthly wage of the male workers in the skilled age range ($w^{\textit{Total}}$) can be defined as follows:
@@
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.106106106Strictly, this weight is not required to be in this range. However, all the cases reported in Kitazawa (1924) take the weights within this range. The monthly wage equation is then specified as follows:
@@
To calculate equation 10, I have to conduct a few final tunings in my wage data. First, I fixed the deviations in the timing of the manufacturing censuses and the THBS observations. Let $w^{\textit{Regular}}_{j}$ be the monthly regular wage calculated using wage statistics measured in year $j\in\{1918,1919\}$ as follows:
@@
where $w^{\textit{Daily}}_{j}$ is the average daily wage of male factory workers in the skilled age range and $l_{j}$ is the average annual working days (Online Appendix B.6.1). The THBS were conducted from November 1918 to January 1920, and my THBS sample was observed from December 1918 to December 1919 (Section 3). However, $w^{\textit{Regular}}_{j}$ is calculated using the wage statistics measured in the manufacturing census conducted in December of 1918 and 1919. This means that if I use $w^{\textit{Regular}}_{1919}$ alone for equation 10, it shall provide upward-biased wage due to the deviation between the timing of the THBS observations and of the manufacturing census. To cope with this issue, I use the weighted average between $w^{\textit{Regular}}_{1918}$ and $w^{\textit{Regular}}_{1919}$ to yield monthly wage series as follows:
@@
where $m\in\{0,1,2,...,12\}$ corresponds to the year-month cells from December 1918 to December 1919. This way of weighting assumes that the average daily wage of machinery workers had increased during this period. This assumption is plausible given the historical fact that the average wage in the manufacturing industry indeed showed a secular increasing trend around 1920 in Tokyo City (Ohkawa et al. 1967, p. 255).
Secondly, I use the number of observations in each year-month cell as the frequency weight. Let $n_{m}$ be the number of households measured in year-month cell $m$. Then, the average monthly wage can be defined as:
@@
Panel A in Table B.4 summarizes all the components in equation 14. In column 1 of Panel B in Table B.4, I calculate the average monthly wage under the share of the ancillary wage for the factories in Tsukishima ($\vartheta$). The estimate is $56$ yen, which is similar to the average and median of monthly earnings in the THBS heads ($59$ and $56$ yen, respectively; see Panel C of Table 2). Therefore, I can conclude that the deviation from the population mean ($56$ yen) is not too large to infer the mean household behaviors of the skilled workers in the machinery factories in Tsukishima. Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $3$ yen difference from the average monthly income is roughly one day’s earnings. This is consistent with the findings in my descriptive analyses that show the similarities in the distribution of family structure, average household size, and share of ancillary wage between the THBS households and Tsukishima households (Section 3; Online Appendix B.6.2).
Column 2 uses the share of the ancillary wage for the entire Tokyo City ($\vartheta=0.09$) to calculate the average monthly wage. The estimate is $54.5$ yen, which is slightly smaller than that in Column 1, reflecting the difference in the weights for the ancillary wages. If I see this figure as an unbiased estimate for the entire city, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in this deviation (Panel B in Table B.3).
@@ B.7 Measurement Errors in the Unadjusted Panel Dataset
Note: Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Unadjusted series defines the days from the 1st to the 15th of each month as the first half, whereas the days from the 16th to the end of each month as the second half. The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half. Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half”.
As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday. The shifted series can deal with the measurement errors from the timing of payday but can not adjust the difference in the payday timing and expenses. In other words, the shifted series ignores the households’ decision on the consumption schedule in each semi-month. The adjusted series can deal with measurement errors and take the timing of the household expenses into account.
Figures B.10bB.10dB.10f illustrate the average semi-monthly income minus expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively.
@@
Figure B.10 explains how different ways of aggregation influence the measured time-series in income and expenditure. Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from raw, shifted, and adjusted panel datasets, respectively.
@@
The unadjusted (calender-month) series defines the days from the 1st to the 15th of each month as the first half and the days from the 16th to the end of each month as the second half. This series shall contain substantial measurement errors due to the mismatches in the income and expenditure data explained in Section 3.3. The time-series plots of average income in Figure B.10a have substantially large variances and fluctuate over time. This is because some semi-monthly income is assigned to the wrong semi-month cells. In these wrong cells, the measured incomes shall take unnaturally greater values than usual. Moreover, the time-series plots of average expenditure do not chase those of the income. For example, the average expenditure in the second half of June greatly differs from the average income in the same cell. Consequently, the average net income (i.e., income minus expenditure) in each semi-month is relatively flat and takes a few odd cells with large positive values (Figure B.10b).
@@
The shifted series defines the days from the 2nd to the 16th of each month as the first half and the days from the 17th to the 1st of the following month as the second half. Although the shifted series can partly deal with the measurement errors in income, it does not adjust the difference in the payday timing and expenses. Figure B.10c indeed gets a much smaller variance in the average incomes over time because all the paydays are now fixed to assign the suitable semi-month cells. However, the expenditure series still does not chase the income series well because the payday is placed on the latter edge in each semi-month cell. In other words, the average expenditure in each semi-month reflects a large part of the consumption based on the income of the previous semi-month cell. Since the average incomes take a better series than Figure B.10a, the average net income offers better time-series plots (Figure B.10d). However, it still has a few cells with large negative and large values due to the deviations of expenditure from income.
@@
The adjusted series uses the days from the 12th to 26th of each month as the “first half” and those from the 27th to 11th of the following month as the “second half”. By shifting the entire semi-month cells, this adjusted series can deal with the measurement errors from the payday shifts and fix the miss-matching in the timing of income and expenditure. In other words, each payday is assigned to a suitable semi-month cell, and each semi-month cell reflects the household’s decision on consumption given the latest income. Figure B.10e shows similar time-series plots of average income in Figure B.10c, in which has a corrected semi-monthly average income series only. Notably, the expenditure series are now chasing the income series in Figure B.10e because each semi-month cell’s expenditure is based on the income in the same cell. Accordingly, Figure B.10f now shows a smoother series in the net income. It shows a clear increasing trend in the net income and has a smaller variance over time.
@@
These illustrations support the evidence that the adjusted semi-monthly datasets can offer cleaner within variations in income and expenditure because they can fix the measurement errors and can reflect the households’ consumption behaviors based on their latest incomes measured in the same semi-month cell.
@@ B.8 Distribution of the Idiosyncratic Shocks
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.107107107I used the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the denser panels include a set of consecutive household-date cells with no income. These cells can not be used to predict the idiosyncratic shocks as the residuals economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
@@ B.9 Unit Root and Autocorrelations
The results of the Fisher-type panel unit root tests based on augmented Dickey-Fuller (ADF) tests are reported in this table. The null hypothesis is that all the panels contain unit roots, whereas the alternative hypothesis is that at least one panel is stationary. In all the specifications, the process under the null hypothesis is assumed to be a random walk with drift. The demeaned data are used to address the effect of cross-sectional dependence. The number of lagged differences in the ADF regression equation reported is set as one because including two or more two lags reduces a large part of the panel units available for computing the test statistics due to the unbalancedness of the panels. I confirmed that including two lags in the semi-monthly panels does not change the results. Including two lags in the adjusted monthly panels and three lags in the semi-monthly panels is impossible. See Choi (2001) for the details of the tests. These tests need both the first-differenced dependent variable and the lagged dependent variable. Therefore, the units with less than four time-series observations are excluded in the regressions.
@@ B.10 Risk Preferences
The potential influences of heterogeneous risk preference on standard risk-sharing regressions are discussed in a few works (Schulhofer-Wohl 2011; Mozzoco and Saini 2012). In my empirical setting, the heterogeneity in the risk preference is assumed to be removed by using the household fixed-effect. When the risk preferences are difference over households, say $\sigma=\sigma_{i}$, equation 26 becomes:
@@
Schulhofer-Wohl (2011) shows that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised under the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock can not be removed using the time-fixed effect in the risk-sharing regressions.108108108Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that can not be applicable to my THBS dataset.109109109The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
@@
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.110110110This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there’s a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
Since the THBS has no experimental questions, it is impossible to apply the same test. Given this limitation that the direct test using experimental risk-tolerance question is unavailable, I compare the correlations between the heads’ income and aggregate consumption over sub-groups to get insights into the potential difference in the heads’ attitudes toward risks. For household $i$ and semi-month $t$, the specification is characterized as follows:
@@
$d_{i}$ is an indicator variable that takes one if household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I consider the case that the heads who worked in the large-scale factory had different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factory might have had a lower risk tolerance than those in the small- and medium-scale enterprises. 13 heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I assume that the heads who worked in the smithing and the non-machinery factories had different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.111111111Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
Table B.6 presents the results. In Column (1), the estimated coefficient on the interaction term is close to zero and statistically insignificant. Thus, the sensitivities to aggregate consumption are similar across different scales of the factories. Column (2) further shows that the estimated coefficient on the interaction term is not statistically significant under an alternative proxy of risk tolerance. It suggests that the heads worked in the smithing and non-machinery factories show similar responses to the aggregate shock. Since the aggregate consumption is the cross-sectional sample average, the aggregate variable may contain noise in the period with the small number of cross-sectional observations. To deal with this potential issue, I exclude October and November 1919 in Columns (3) and (4). As shown, the estimated coefficients on the interaction terms are largely unchanged in both columns.
This finding is consistent with the fact that these heads have similar characteristics in nature because the THBS focused on skilled factory workers in a representative manufacturing area, Tsukishima. Therefore, it is plausible to see that the THBS households had more or less similar attitudes toward the risks.
@@
(1) Full (2) Full Jan. to Sept. Jan. to Sept.
@@
Notes: This table shows the results of equation 19 for the semi-monthly panel dataset. ‘Aggregate Consumption’ indicates the cross-sectional sample average of the total expenditures. ‘Large Scale Factory’ indicates an indicator variable that takes one if the head worked in the Ishikawajima Shipyard or governmental factory. ‘Smithing/Non-machinery’ indicates an indicator variable that takes one if the head worked in the smithing, food, or paper-making factories. All the regressions include the family size variable and household fixed-effect.
Columns 1–2 show the results for the entire sample period. Columns 2–3 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).

### v10

@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
1. Average daily wage ($w^{Daily}_{j}$): The average daily wage of male factory workers in the skilled age range (aged 30–40) in the machinery sector is calculated using two official reports of the manufacturing censuses documented in the ASCT and LST. Details in the calculation steps are summarized in Online Appendix B.6.1.
2. Average annual working days ($l_{j}$): The ASCT documents the number of factories and average annual working days for each machinery sub-sector in two factory-size bins. I use the number of factories in each machinery sub-sector and factory size bin as the weights to calculate the weighted average of the annual working days in the entire machinery sector.
@@
Sources: Data used to calculate the average daily wage are from the Tokyo City Office (1921, pp. 726-757); Tokyo City Office (1922, pp. 732-733); Tokyo City Statistics Division (1926a, pp. 16; 244–245). Data used to calculate the average annual working days are from the Tokyo City Office (1921, pp. 726–741) and Tokyo City Office (1922, pp. 716–725). Data used to calculate the average share of monthly ancillary wage are from Kitazawa (1924, pp. 20; 58–62) and the Ministry of Agriculture and Commerce (1921, pp. 613–629; 674).
Following the institution described in Online Appendix B.6.2, the monthly wage of the male workers in the skilled age range ($w^{\textit{Total}}$) is defined as follows:
@@
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.102102102Strictly, this weight may not be required to be in this range. I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one. The monthly wage equation is then specified as follows:
@@
To calculate equation 10, I have to conduct a few final tunings in my wage data. First, I fix the deviations in the timing of the manufacturing censuses and the THBS observations. Let $w^{\textit{Regular}}_{j}$ be the monthly regular wage calculated using wage statistics measured in year $j\in\{1918,1919\}$ as follows:
@@
where $w^{\textit{Daily}}_{j}$ is the average daily wage of male factory workers in the skilled age range and $l_{j}$ is the average annual working days (Online Appendix B.6.1). The THBS were conducted from November 1918 to January 1920, and my THBS sample was observed from December 1918 to December 1919 (Section 3). However, $w^{\textit{Regular}}_{j}$ is calculated using the wage statistics measured in the manufacturing census conducted in December of 1918 and 1919. This means that if I used $w^{\textit{Regular}}_{1919}$ alone for equation 10, it shall provide upward-biased wage due to the deviation between the timing of the THBS observations and of the manufacturing census. To manage this issue, I use the weighted average between $w^{\textit{Regular}}_{1918}$ and $w^{\textit{Regular}}_{1919}$ to yield monthly wage series as follows:
@@
where $m\in\{0,1,2,...,12\}$ corresponds to the year-month cells from December 1918 to December 1919. This way of weighting assumes that the average daily wage of machinery workers had increased during this period. This assumption is plausible given the historical fact that the average wage in the manufacturing industry indeed showed a secular increasing trend circa 1920 in Tokyo City (Ohkawa et al. 1967, p. 255).
Second, I use the number of observations in each year-month cell as the frequency weight. Let $n_{m}$ be the number of households measured in year-month cell $m$. Then, the average monthly wage can be defined as:
@@
Panel A in Table B.4 summarizes all the components in equation 14. In Column 1 of Panel B in Table B.4, I calculate the average monthly wage under the share of the ancillary wage for the factories in Tsukishima ($\vartheta$). The estimate is $56$ yen, which is similar to the average and median of monthly earnings in the THBS heads ($59$ and $56$ yen, respectively; see Panel C of Table 2). Therefore, I can conclude that the deviation from the population mean ($56$ yen) is not too large to infer the mean household strategies of the skilled workers in the machinery factories in Tsukishima. Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $3$ yen difference from the average monthly income is approximately one day’s earnings. This is consistent with the findings in my descriptive analyses that show the similarities in the distribution of family structure, average household size, and share of ancillary wage between the THBS households and Tsukishima households (Section 3; Online Appendix B.6.2).
Column 2 uses the share of the ancillary wage for the entire Tokyo City ($\vartheta=0.09$) to calculate the average monthly wage. The estimate is $54.5$ yen, which is slightly smaller than that in Column 1, reflecting the difference in the weights for the ancillary wages. If I view this figure as an unbiased estimate for the entire Tokyo City, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in this deviation (Panel B in Table B.3).
@@ B.7 Measurement Errors in the Unadjusted Panel Dataset
Note: Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Unadjusted series defines the days from the 1st to the 15th of each month as the first half, whereas the days from the 16th to the end of each month as the second half. The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half. Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half.”
As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday. The shifted series can manage the measurement errors from the timing of payday but cannot adjust the difference in the payday timing and expenses. In other words, the shifted series ignores the households’ decision on the consumption schedule in each semi-month. The adjusted series reflecting the exact timing of the household expenses can manage measurement errors. Figures B.10bB.10dB.10f illustrate the average semi-monthly income minus expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively.
@@
Figure B.10 explains how different methods of aggregation influence the time-series in income and expenditure. Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from unadjasted, shifted, and adjusted panel datasets, respectively.
@@
The unadjusted (calender-month) series defines the days from the 1st to the 15th of each month as the first half and the days from the 16th to the end of each month as the second half. This series shall contain substantial measurement errors due to the mismatches in the income and expenditure data explained in Section 3.3. The time-series plots of average income in Figure B.10a have a substantially large variance and fluctuate over time. This is because some semi-monthly income is assigned to the wrong semi-month cells. In these wrong cells, the measured incomes take unnaturally greater values than usual. Moreover, the time-series plots of average expenditure do not chase those of average income. For example, the average expenditure in the second half of June greatly differs from the average income in the same cell. Consequently, the average net income (i.e., income minus expenditure) in each semi-month takes unnaturally large positive values in a few cells (Figure B.10b).
@@
The shifted series defines the days from the 2nd to the 16th of each month as the first half and the days from the 17th to the 1st of the following month as the second half. Although the shifted series can partly manage the measurement errors in income, it does not adjust the difference in the timing of paydays and expenses. Figure B.10c obtains a much smaller variance in the average incomes over time because all the paydays are now fixed to assign suitable semi-month cells. However, the expenditure series still does not chase the income series well because the payday is placed on the latter edge in each semi-month cell. In other words, the average expenditure in each semi-month reflects a large part of the consumption based on the income of the previous semi-month cell. Since the average incomes take a better series than Figure B.10a, the average net income offers better time-series plots (Figure B.10d). However, it still has a few cells with large negative and large values due to the deviations of expenditure from income.
@@
The adjusted series uses the days from the 12th to 26th of each month as the “first half” and those from the 27th to 11th of the following month as the “second half.” By shifting the entire semi-month cells, this adjusted series can manage the measurement errors from the payday shifts and can fix the miss-matching in the timing of income and expenditure. In other words, each payday is assigned to a suitable semi-month cell, and each semi-month cell reflects the household’s decision on consumption in response to the latest income. Figure B.10e shows similar time-series plots of average income in Figure B.10c, in which has a corrected semi-monthly average income series only. Notably, the expenditure series are now chasing the income series in Figure B.10e because each semi-month cell’s expenditure is based on the income in the same cell. Accordingly, Figure B.10f now shows a smoother series with a smaller variance in the net income.
@@
These illustrations support the evidence that the adjusted semi-monthly datasets can offer cleaner within variations in income and expenditure because they can fix the measurement errors and reflect the households’ consumption strategies based on their latest incomes measured in the same semi-month cell.
@@ B.8 Distribution of the Idiosyncratic Shocks
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.103103103I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the daily panels include a set of consecutive household-date cells with no income. These cells cannot be used to predict the idiosyncratic shocks as the residuals are economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
@@ B.9 Unit Root and Autocorrelations
Notes: The results of the Fisher-type panel unit root tests based on augmented Dickey-Fuller (ADF) tests are reported in this table. The null hypothesis is that all the panels contain unit roots, whereas the alternative hypothesis is that at least one panel is stationary. In all the specifications, the process under the null hypothesis is assumed to be a random walk with drift. The demeaned data are used to address the effect of cross-sectional dependence. The number of lagged differences in the ADF regression equation reported is set as one because including two or more two lags reduces a large part of the panel units available for computing the test statistics due to the lack of balance of the panels. I confirm that including two lags in the semi-monthly panels does not change the results. Including two lags in the adjusted monthly panels and three lags in the semi-monthly panels is impossible. See Choi (2001) for the details of the tests. These tests need both the first-differenced dependent variable and the lagged dependent variable. Therefore, the units with less than four time-series observations are excluded in the regressions.
@@ B.10 Risk Preferences
The potential influences of heterogeneous risk preference on standard risk-sharing regressions are discussed in a few works (Schulhofer-Wohl 2011; Mozzoco and Saini 2012). In my empirical setting, the heterogeneity in the risk preference is assumed to be removed by using the household fixed-effect. When the risk preferences are differenced over households like $\sigma=\sigma_{i}$, equation 26 becomes:
@@
Schulhofer-Wohl (2011) shows a specific case (utility function) that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.104104104Specifically, Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.105105105The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
@@
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.106106106This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
Since the THBS has no experimental questions, it is impossible to apply the same test. Given this limitation—that the direct test using experimental risk-tolerance question is unavailable—I compare the correlations between the heads’ income and aggregate consumption over sub-groups to obtain insights into the potential difference in the heads’ attitudes toward risks. For household $i$ and semi-month $t$, the specification is characterized as follows:
@@
$d_{i}$ is an indicator variable that takes one if a household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I test the hypothesis that the heads who work in the large-scale factories have different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factories might have had a lower risk tolerance than those in the small- and medium-scale enterprises. Thirteen heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.107107107Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
Table B.6 presents the results. In Column (1), the estimated coefficient on the interaction term is close to zero and statistically insignificant. Thus, the sensitivities to aggregate consumption are similar across different scales of factories. Column (2) further shows that the estimated coefficient on the interaction term is not statistically significant under an alternative proxy of risk tolerance. It suggests that the heads who work in the smithing and non-machinery factories show similar responses to the aggregate shock. Since the aggregate consumption is the cross-sectional sample average, the aggregate variable may contain noise in the period with the small number of cross-sectional observations. To manage this potential issue, I exclude October and November 1919 in Columns (3) and (4). As shown, the estimated coefficients on the interaction terms are largely unchanged in both columns.
This finding is consistent with the fact that these heads have similar characteristics because the THBS focused on skilled factory workers in a representative manufacturing area, Tsukishima. Therefore, it is plausible that the THBS households had similar attitudes toward the risks.
@@
(1) Full (2) Full (3) Jan. to Sept. (4) Jan. to Sept.
@@
Notes: This table shows the results of equation 19 for the semi-monthly panel dataset. ‘Aggregate Consumption’ indicates the cross-sectional sample average of the total expenditures. ‘Large Scale Factory’ indicates an indicator variable that takes one if the head worked in the Ishikawajima Shipyard or governmental factory. ‘Smithing/Non-machinery’ indicates an indicator variable that takes one if the head worked in the smithing, food, or paper-making factories. All the regressions include the family size variable and household fixed effect. Columns 1–2 show the results for the entire sample period. Columns 3–4 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).